# Advancing Multimodal Reasoning: From Optimized Cold Start to Staged Reinforcement Learning

### v1

@@ Abstract
Inspired by the remarkable reasoning capabilities of Deepseek-R1 in complex textual tasks, many works attempt to incentivize similar capabilities in Multimodal Large Language Models (MLLMs) by directly applying reinforcement learning (RL). However, they still struggle to activate complex reasoning. In this paper, rather than examining multimodal RL in isolation, we delve into current training pipelines and identify three crucial phenomena: 1) Effective cold start initialization is critical for enhancing MLLM reasoning. Intriguingly, we find that initializing with carefully selected text data alone can lead to performance surpassing many recent multimodal reasoning models, even before multimodal RL. 2) Standard GRPO applied to multimodal RL suffers from gradient stagnation, which degrades training stability and performance. 3) Subsequent text-only RL training, following the multimodal RL phase, further enhances multimodal reasoning. This staged training approach effectively balances perceptual grounding and cognitive reasoning development. By incorporating the above insights and addressing multimodal RL issues, we introduce ReVisual-R1, achieving a new state-of-the-art among open-source 7B MLLMs on challenging benchmarks including MathVerse, MathVision, WeMath, LogicVista, DynaMath, and challenging AIME2024 and AIME2025.
@@ 1 Introduction
Recently, the field of large language models (LLMs) has witnessed significant advancements in complex cognitive reasoning zeng7b ; yan2025learning , notably exemplified by reasoning models like DeepSeek-R1 guo2025deepseek . These models successfully leveraged Reinforcement Learning (RL) to facilitate the self-emergence of intricate reasoning abilities in text-only models. Inspired by this success, a natural extension has been to apply similar RL paradigms to Multimodal Large Language Models (MLLMs) with the goal of incentivizing multimodal reasoning capabilities visionr1 ; meng2025mm ; dong2024insight ; sun2024surf ; xia2025mmed ; peng2025lmm ; su2025openthinkimg ; shen2025satori .
Despite these promising efforts, current methods often struggle to fully unlock complex reasoning within MLLMs. This difficulty suggests that directly transplanting RL techniques from text-only domains may not adequately address the unique dynamics and requirements inherent in multimodal learning and reasoning. Motivated by this gap, our work comprehensively studies the training pipelines of the multimodal reasoning models. Through this investigation, we identify three crucial phenomena that significantly influence the efficacy of multimodal training:
@@
First, we observe that sufficient cold start initialization is indispensable for effectively cultivating the reasoning ability of MLLMs. Conventional cold-start phases for MLLMs, often relying on simplistic visual and textual pre-training corpora, appear to insufficiently prepare models for the demands of complex problem-solving, a challenge evident in the reasoning limitations of various existing models r1-onevision ; visionr1 ; wang2025vl ; openvlthinker ; chen2025r1v . This initial deficit critically hinders the subsequent RL stages from eliciting sophisticated, self-critical reasoning patterns. To unlock deeper deliberative reasoning in MLLMs, an enriched cold-start initialization is therefore not merely beneficial, but indispensable. Specifically, initializing with carefully selected text data that instills foundational reflective capabilities and the capacity for extended Chain-of-Thought (CoT) reasoning proves to be a powerful strategy. Intriguingly, such targeted textual initialization allows our model to surpass the multimodal reasoning performance of many recent multimodal reasoning models.
@@
Second, we identify that the standard Group Relative Policy Optimization (GRPO) algorithm shao2024deepseekmath , commonly applied for multimodal RL, suffers from a gradient stagnation problem. This issue significantly degrades both the training stability and the ultimate performance of the multimodal RL phase. To address this fundamental limitation and improve the efficacy of multimodal RL, we propose Prioritized Advantage Distillation (PAD). PAD is designed to mitigate gradient stagnation by strategically filtering out zero-advantage samples and re-weighting informative trajectories, thereby focusing the learning process on more impactful data and improving training stability.
Third, we discover that conducting further post-training using text RL after the multimodal RL training phase further enhances multimodal reasoning ability. This observation highlights the complex interplay between visual grounding and linguistic fluency in achieving superior multimodal reasoning. Based on these three key findings, we devise a three-stage curriculum that effectively balances visual and textual competencies for multimodal reasoning. This curriculum comprises: (1) a textual cold-start to explicitly instill complex reasoning templates; (2) a multimodal RL stage to connect linguistic reasoning with visual perception; and (3) a text-only RL refinement stage aims to restore linguistic fluency and refine reasoning expression without eroding the newly gained visual grounding skills, effectively consolidating the multimodal reasoning capabilities.
@@
To this end, we introduce ReVisual-R1, the first 7B-parameter open-source MLLM designed with this staged curriculum. As shown in Figure 1, ReVisual-R1 presents strong performance in challenging multimodal reasoning tasks, while simultaneously preserving strong general-purpose textual skills. Extensive experiments on a suite of challenging benchmarks, including MathVerse zhang2024mathverse , MathVision wang2024measuring , MathVista lu2023mathvista , DynaMath zou2025dynamathdynamicvisualbenchmark , WeMath qiao2024wemathdoeslargemultimodal , and LogicVista xiao2024logicvistamultimodalllmlogical , as well as the AIME24/25 li2024numinamath , GPQA rein2024gpqa , MATH-500 dataset_math benchmark, confirm that ReVisual-R1 significantly outperforms much larger public models.
@@
  • We systematically investigate MLLM cold-start initialization, revealing the insufficiency of existing multimodal pre-training datasets and demonstrating that a text-centric, high-difficulty cold-start phase is crucial for unlocking complex multimodal reasoning capabilities.
  • We identify and address the critical issue of gradient stagnation in GRPO for multimodal RL by proposing Prioritized Advantage Distillation (PAD), a novel technique that ensures more stable training and sample-efficient learning for MLLMs.
@@ 2.1 Multimodal Reasoning Formulation
where $\mathcal{D}$ represents the distribution of multimodal reasoning tasks. Similar to Deepseek R1 guo2025deepseek , in our work, we mainly use rule-based reward, $r(x,y)=1$ if $y$ is correct, otherwise $r(x,y)=0$.
@@ 2.2 Group Relative Policy Optimization
where $\epsilon$ is a small constant for numerical stability. This relative advantage is then used within a clipped surrogate objective function.
Here, $r(x,y_{i})$ represents the reward for response $y_{i}$ to input $x$. This advantage function measures how much better a specific response is compared to the average performance within its group, normalized by the group’s reward variance. By using group-specific advantages, GRPO encourages the model to improve responses within each group while maintaining diversity across groups.
3 GRAMMAR: Generalized Multimodal Reasoning Dataset
In this section, we first show an intriguing finding involving the multimodal reasoners in Section 3.1, paving the way for the strategy of our data curation pipeline in Section 3.2.
@@ 3.1 Preliminary Study of Cold Start
FigureQA figureqa-samira-iclr-2019 100K Super-CLEVR superclevr-li-cvpr-2023 30K Big-Math-RL bigmath-alon-arxiv-2025 251K GAIR_LIMO ye2025limoreasoning 0.8K
MAVIS mavis-zhang-iclr-2025 218K TabMWP tabmwp-lu-iclr-2023 38K Big-Math-RL-U 35K s1K-1.1 muennighoff2025s1simpletesttimescaling 1K
GeoQA geoqa-chen-acl-2021 5K UniGeo unigeo-chen-emnlp-2022 16K OpenThoughts openthoughts 114K OpenMathR moshkov2025aimo2 3,200K
Geometry3K geometry3k-lu-acl-2021 2.1K MultiMath multimath-peng-arxiv-2024 300K DeepMath he2025deepmath103klargescalechallengingdecontaminated 103K OrcaMath orcamath-arindam-arxiv-2024 200K
IconQA iconqa-lu-nips-2021 107K OpenR1-220k openr1 220K NuminaMath-CoT numina_math_datasets 859K
Figure 2: Absolute performance improvement on Qwen2.5-VL-7B-Instruct across textual and multimodal reasoning tasks. The red and purple dashed lines represent the average absolute gains of VisionR1/R1-One-Vision and DeepMath/OpenR1-Math over the baseline, respectively, across four reasoning tasks.
Currently, the leading text and multimodal reasoning models in the community guo2025seedvl ; guo2025deepseek primarily rely on cold-start training with extensive self-critique data or large-scale multimodal reinforcement learning based on verifiable rewards guo2025deepseek . While cold-start training is essential for improving the reasoning capabilities of models guo2025deepseek , its effect on the reasoning abilities of multimodal models remains underexplored. To investigate this, we collect two open-source cold-start multimodal datasets, Vision-R1 visionr1 and R1-One-Vision r1-onevision , along with two cold-start textual datasets, DeepMath he2025deepmath103klargescalechallengingdecontaminated and OpenR1-Math openr1 . Then we randomly sample 40,000 instances from these datasets to fine-tune Qwen2.5-VL-7B-Instruct bai2025qwen2 . The fine-tuned models are subsequently evaluated on multimodal reasoning benchmarks (MathVerse and MathVision) as well as text reasoning benchmarks (AIME24 and Math500). The experimental outcomes and average performance enhancements from the multimodal and textual cold-start datasets are illustrated in Figure 2.
To further investigate this observation, we perform an analysis using a subset of 100 examples sampled from the Vision-R1 visionr1 and DeepMath he2025deepmath103klargescalechallengingdecontaminated datasets. Specifically, we analyze the response lengths and pass rates of the doubao-1.5-thinking-pro-vision model seed2025seed1 on these samples. Responses to textual prompts from DeepMath averaged 8,207.76 tokens, which is substantially longer than the 821.48 tokens generated in response to multimodal prompts from Vision-R1. Moreover, the pass rate for Vision-R1 is 96.00%, whereas DeepMath achieve a pass rate of only 75.0%. These findings further indicate that current multimodal cold start datasets may lack sufficient complexity to inspire advanced reasoning capabilities of reasoning models. It indicates that existing multimodal cold start datasets in the community may not be sufficiently challenging to enhance the complex reasoning capabilities and generalization of policy models, i.e., they may not provide effective initial strategies during cold start training. Therefore, it is desirable to develop a data curation pipeline for multimodal policy model to improve the generalization capabilities.
3.2 Data Curation
Informed by Section 3.1 regarding the variability in open-source reasoning data lu2022iconqanewbenchmarkabstract , chen2022geoqageometricquestionanswering , zhang2024mavis , lindström2022clevrmathdatasetcompositionallanguage , chang2022mapqadatasetquestionanswering , we develop GRAMMAR, a new dataset designed to enhance the generalization of reasoning capabilities in multimodal models. GRAMMAR comprises 47k diverse textual thought samples with explicit reasoning paths, augmented by 31k complex textual examples and 21k multimodal questions with ground truth annotations suitable for rule-based reinforcement learning.
@@ 3.2 Data Curation
The construction of GRAMMAR involved a multi-stage curation pipeline. We begin by amassing a wide array of open-source reasoning data, spanning various difficulty levels (details in Table 1). This initial collection underwent rule-based filtering to ensure answer verifiability, excluding items like proof problems and those with difficult-to-verify ground truths. Subsequently, Qwen2.5-VL-7B-Instruct was employed for initial pruning of overly simple or complex questions. Qwen2.5-VL-32B-Instruct is then used to assess the remaining samples to classify them into ten difficulty levels. To maximize data diversity and minimize redundancy, we encoded questions using NV-Embedding-V2 lee2024nv , applied HDBSCAN campello2013density for clustering, assigned topics to clusters via Qwen2.5-7B-Instruct, and performed balanced sampling across both topics and difficulty strata.
@@ 4 Staged Reinforcement Optimization (SRO)
Our data investigations (Section 3.1) and the curation of the GRAMMAR dataset highlight the necessity of high-quality, reasoning-focused data for developing advanced MLLM capabilities. Building directly on these data-centric foundations, we introduce Staged Reinforcement Optimization (SRO), a framework designed to systematically cultivate robust reasoning and diverse competencies in MLLMs. SRO achieves this through a sequence of distinct learning phases, each tailored to address specific training challenges and leverage appropriate components of the GRAMMAR dataset. This section details the SRO architecture and its constituent techniques.
@@ 4.1 Stage 1: Multimodal RL
After the cold start training, the SRO framework commences with a dedicated Multimodal Reinforcement Learning (MRL) phase. This initial stage is pivotal for enabling the MLLM to ground textual concepts in visual information and execute cross-modal reasoning, primarily using the multimodal samples from our GRAMMAR dataset. We employ GRPO as the core RL algorithm for this phase. To ensure stable and effective learning, particularly when dealing with complex tasks and potentially sparse rewards common in multimodal settings, we integrate two key enhancements: Prioritized Advantage Distillation (PAD) to improve gradient quality by addressing specific GRPO limitations, and an efficient-length reward function for more controlled and stable response generation.
@@ 4.1.1 Prioritized Advantage Distillation (PAD)
In this paper, we discover a significant challenge when applying GRPO in complex multimodal settings is “Gradient Stagnation”. This phenomenon refers to a reduction in learning efficacy due to a predominance of near-zero advantage estimates, which is particularly acute when dealing with sparse binary rewards. Essentially, if entire groups of generated responses yield uniform rewards (e.g., all correct or all incorrect), the resulting advantage signals become null, leading to zero policy gradients and thereby halting learning for those samples. This issue, also noted in concurrent works wang2025vl ; yu2025dapo , can severely impede training progress. To specifically counteract gradient stagnation and enhance the efficiency of GRPO, we introduce Prioritized Advantage Distillation (PAD). PAD refines the training process by strategically focusing updates on the most informative samples within each batch, namely those exhibiting significant, non-zero advantage signals. This approach optimizes computational resource allocation and promotes more consistent learning. The PAD mechanism, detailed below, operates on each batch after initial advantage estimation:
  • Per-Sequence Advantage Calculation: Compute the absolute advantage $|\tilde{A}_{i}|$ for each sequence $i$ in the original batch $\mathcal{B}$, representing its learning signal magnitude.
  • Effective Sample Filtering: Form an “effective set” $\mathcal{E}$ by selecting sequences $i$ whose absolute advantage $|\tilde{A}_{i}|$ falls within a specified informative range $[T_{low},T_{high}]$. Critically, $T_{low}>0$ filters out stagnant (near-zero advantage) samples, ensuring that candidates for sub-sampling provide potentially useful learning signals.
@@
PAD thus directly counteracts gradient stagnation via a dual mechanism: first, by filtering out stagnant samples, and second, by prioritizing updates using informative, non-zero advantages from the remaining set. This selective optimization of the learning process ensures efficient computational resource allocation towards high-value samples. Consequently, PAD leads to enhanced training stability, improved learning efficiency, and more effective acquisition of complex reasoning skills, particularly in challenging scenarios with sparse or binary rewards.
4.1.2 Efficient-Length Reward Function
While complex reasoning tasks often necessitate extended outputs, excessively long sequences can be suboptimal qu2025survey ; fu2025scaling . Therefore, in this paper, besides the primary reward signal for task accuracy, we introduce an efficient-length reward to regulate the verbosity of generated responses. Specifically, let $L_{y}$ be the token length of the generated sequence and $L_{\text{budget}}$ be a pre-defined target length budget. A raw reward score, $R_{\text{raw}}$, is first computed as a linear function of the deviation:
$R_{\text{raw}}=\alpha(L_{\text{budget}}-L_{y})+\delta$ (5)
where $\alpha$ is a positive scaling factor controlling the penalty magnitude for length deviation, and $\delta$ is a baseline reward (e.g., 0.5) assigned when $L_{y}=L_{\text{budget}}$. The final Efficient-Length Reward, $R_{\text{len}}$, is obtained by clipping $R_{\text{raw}}$ to the range $[0,1]$:
$R_{\text{len}}(L_{y},L_{\text{budget}},\alpha,\delta)=\max(0.0,\min(1.0,R_{\text{raw}}))$ (6)
In our experiments, we typically use a small $\alpha$ (e.g., 0.005) to ensure a gentle slope for the reward adjustment. This formulation provides a continuous and bounded reward signal. Specifically, a reward of $\delta$ is assigned if the generated length $L_{y}$ matches the budget $L_{\text{budget}}$. The reward proportionally increases (up to a maximum of 1.0) as $L_{y}$ becomes shorter than $L_{\text{budget}}$, thereby encouraging conciseness. This Efficient-Length Reward guides the model towards producing responses that are not only accurate but also parsimonious, without prematurely curtailing potentially valuable, longer reasoning paths, thus fostering more robust and efficient learning of complex multimodal reasoning.

### v2

# Revisual-R1: Advancing Multimodal Reasoning From Optimized Cold Start to Staged Reinforcement Learning

@@ Abstract
Inspired by the remarkable reasoning capabilities of Deepseek-R1 in complex textual tasks, many works attempt to incentivize similar capabilities in Multimodal Large Language Models (MLLMs). However, they still struggle to activate complex reasoning. In this paper, rather than examining multimodal RL in isolation, we delve into current training pipelines and identify three crucial phenomena: 1) Effective cold start initialization is critical for enhancing MLLM reasoning. Intriguingly, we observe that initializing with carefully selected text data alone can lead to performance surpassing many recent multimodal reasoning models, even before multimodal RL. 2) Standard GRPO applied to multimodal RL suffers from gradient stagnation, which degrades both training stability and performance. 3) A final text-only RL tuning stage, conducted after the multimodal RL phase, is effective at further sharpening multimodal reasoning capabilities. Capitalizing on these insights, we introduce ReVisual-R1, achieving a new state-of-the-art among open-source 3B and 7B MLLMs on challenging benchmarks, including MathVerse, MathVision, WeMath, LogicVista, DynaMath, and challenging AIME2024 and AIME2025. This paradigm demonstrates robust scalability, proving effective at both 7B and 3B scales. Our findings chart a new course for training powerful multimodal reasoners, demonstrating that a carefully orchestrated, multi-stage strategy is key to unlocking their full potential.
@@ 1 Introduction
Recently, the field of large language models (LLMs) has witnessed significant advancements in complex cognitive reasoning (Zeng et al., ; Yan et al., 2025; Zhang et al., 2025a), notably exemplified by reasoning models like DeepSeek-R1 (Guo et al., 2025a). These models successfully leveraged Reinforcement Learning (RL) to facilitate the self-emergence of intricate reasoning abilities in text-only models. A natural and ambitious next step is to extend this RL paradigm to Multimodal Large Language Models (MLLMs), with the goal of unlocking a similar leap in multimodal cognitive abilities (Huang et al., 2025b; Meng et al., 2025b; Xia et al., ; Peng et al., 2025).
However, this ambition has been met with a formidable challenge. The direct application of text-centric RL techniques to MLLMs has yielded diminishing returns, suggesting that the path to multimodal reasoning is not a simple matter of architectural extension. A fundamental question arises: how does one cultivate the abstract linguistic skill while simultaneously grounding it in the continuous, high-dimensional space of visual perception? Naive co-training often results in a compromise, where neither modality’s potential is fully realized. Motivated by this critical impasse, we undertake a rigorous dissection of the MLLM training pipeline. Our investigation uncovers three foundational insights that, when addressed in concert, chart a new and more effective course for developing powerful multimodal reasoners.
@@
First, we observe that sufficient cold start initialization is indispensable for effectively cultivating the reasoning ability of MLLMs. Conventional cold-start phases for MLLMs often rely on simplistic visual and textual pre-training corpora (Yang et al., 2025; Huang et al., 2025b; Wang et al., 2025; Deng et al., 2025; Chen et al., 2025b; Feng et al., 2025). This initial deficit critically hinders the subsequent RL stages from eliciting sophisticated, self-critical reasoning patterns. To unlock deeper deliberative reasoning in MLLMs, an enriched cold-start initialization is therefore not merely beneficial but indispensable. Specifically, initializing with carefully selected text data that instills foundational reflective capabilities and the capacity for extended Chain-of-Thought (CoT) reasoning proves to be a powerful strategy. Intriguingly, such targeted textual initialization allows our model to surpass the multimodal reasoning performance of many recent multimodal reasoning models.
@@
Second, we identify that the standard Group Relative Policy Optimization (GRPO) algorithm (Shao et al., 2024), commonly applied for multimodal RL, suffers from a gradient stagnation problem. This issue significantly degrades both the training stability and the ultimate performance of the multimodal RL phase. To address this fundamental limitation and improve the efficacy of multimodal RL, we propose Prioritized Advantage Distillation (PAD). PAD is designed to mitigate gradient stagnation by strategically filtering out zero-advantage samples and re-weighting informative trajectories, thereby focusing the learning process on more impactful data and improving training stability.
Third, we discover that conducting further post-training using text RL after the multimodal RL training phase can further enhance multimodal reasoning ability. This stage acts as a polishing step, sharpening the model’s linguistic expression and logical consistency without eroding the newly acquired visual grounding. It consolidates the model’s capabilities, leading to a synergistic whole greater than the sum of its parts.
@@
Synthesizing these insights, we propose a principled, three-stage training curriculum that strategically sequences these learning phases: (1) a text-centric cold-start to forge a powerful reasoning engine, (2) a multimodal RL phase with PAD to ground this engine in vision, and (3) a text-only RL refinement to consolidate and polish the integrated skills. The culmination of this methodology is ReVisual-R1, the first 7B-parameter open-source MLLM architected around this curriculum. Extensive experiments on a suite of challenging benchmarks, including MathVerse (Zhang et al., 2024b), MathVision (Wang et al., 2024a), MathVista (Lu et al., 2023a), DynaMath (Zou et al., 2025), WeMath (Qiao et al., 2024b), and LogicVista (Xiao et al., 2024), as well as the AIME24/25 (Li et al., 2024), GPQA (Rein et al., 2024), MATH-500 (Hendrycks et al., 2021) benchmark, confirm that ReVisual-R1 significantly outperforms much larger public models. We further validate the scalability of our framework by demonstrating its effectiveness at the 3B model scale.
@@
  • We challenge the conventional MLLM training paradigm by demonstrating that a text-centric, high-difficulty cold-start is the crucial, and previously overlooked, foundation for unlocking elite multimodal reasoning.
  • We identify the fundamental problem of gradient stagnation in multimodal RL and propose Prioritized Advantage Distillation (PAD), a novel and effective solution that stabilizes training and enhances sample efficiency.
@@ 2.1 Multimodal Reasoning Formulation
where $\mathcal{D}$ represents the distribution of multimodal reasoning tasks. Similar to Deepseek R1 (Guo et al., 2025a), we mainly use rule-based reward, $r(x,y)=1$ if $y$ is correct, otherwise $r(x,y)=0$.
@@ 2.2 Group Relative Policy Optimization
where $\epsilon$ is a small constant for numerical stability. This relative advantage is then used within a clipped surrogate objective function. $r(x,y_{i})$ represents the reward for response $y_{i}$ to input $x$. This advantage function measures how much better a specific response is compared to the average performance within its group, normalized by the group’s reward variance.
3 Optimized Cold Start for Multimodal Reasoning
Figure 2: (Top): the overview of our proposed ReVisual-R1 framework. After collecting and curating data, ReVisual-R1 contains cold start and staged reinforcement learning. (Bottom): the process of our proposed prioritized advantage distillation (PAD) for multimodal reinforcement learning.
In this section, we first show an intriguing finding in the cold start of multimodal reasoning in Section 3.1, paving the way for the strategy of our data curation pipeline in Section 3.2. As shown in Figure 2, our Revisual-R1 includes a cold start stage followed by staged reinforcement learning.
@@ 3.1 Preliminary Study of Cold Start
Figure 3: Absolute performance improvement on Qwen2.5-VL-7B-Instruct across textual and multimodal reasoning tasks. The purple and red dashed lines represent the average absolute gains of VisionR1/R1-One-Vision and DeepMath/OpenR1-Math over the baseline, respectively, across four reasoning tasks.
To investigate the effectiveness of cold-start training (Guo et al., 2025a), we first collect two open-source cold-start multimodal datasets, Vision-R1 (Huang et al., 2025b) and R1-One-Vision (Yang et al., 2025), along with two cold-start textual datasets, DeepMath (He et al., 2025) and OpenR1-Math (Face, 2025). Then we randomly sample 40,000 instances from these datasets to fine-tune Qwen2.5-VL-7B-Instruct (Bai et al., 2025). The fine-tuned models are subsequently evaluated on multimodal reasoning benchmarks (MathVerse and MathVision) as well as textual reasoning benchmarks (AIME24 and Math500). The experimental outcomes and average performance enhancements from the multimodal and textual cold-start datasets are illustrated in Figure 3.
To further investigate this observation, we perform an analysis using a subset of 100 examples sampled from the Vision-R1 (Huang et al., 2025b) and DeepMath (He et al., 2025) datasets. Specifically, we analyze the response lengths and pass rates of the doubao-1.5-thinking-pro-vision model (Seed et al., 2025) on these samples. Responses to textual prompts from DeepMath averaged 8,207.76 tokens, which is substantially longer than the 821.48 tokens generated in response to multimodal prompts from Vision-R1. Moreover, the pass rate for Vision-R1 is 96.00%, whereas DeepMath achieves a pass rate of only 75.0%. These findings further indicate that current multimodal cold start datasets may lack sufficient complexity to inspire advanced reasoning capabilities of reasoning models. Therefore, in this paper, we adopt textual-only data for the cold start stage.
3.2 GRAMMAR: Generalized Multimodal Reasoning Dataset
@@
FigureQA (Kahou et al., 2018) 100K Super-CLEVR (Li et al., 2023) 30K Big-Math-RL (Albalak et al., 2025) 251K GAIR_LIMO (Ye et al., 2025) 0.8K
MAVIS (Zhang et al., 2025b) 218K TabMWP (Lu et al., 2023b) 38K Big-Math-RL-U 35K s1K-1.1 (Muennighoff et al., 2025) 1K
GeoQA (Chen et al., 2021) 5K UniGeo (Chen et al., 2022) 16K OpenThoughts (Team, 2025) 114K OpenMathR (Moshkov et al., 2025) 3,200K
Geometry3K (Lu et al., 2021a) 2.1K MultiMath (Peng et al., 2024) 300K DeepMath (He et al., 2025) 103K OrcaMath (Mitra et al., 2024a) 200K
IconQA (Lu et al., 2021b) 107K OpenR1-220k (Face, 2025) 220K NuminaMath-CoT (LI et al., 2024) 859K
Informed by Section 3.1, in this paper, we develop the Generalized Multimodal Reasoning Dataset (GRAMMAR) to enhance the generalization of reasoning capabilities in multimodal models. Specifically, the GRAMMAR dataset consists of two components. For the cold-start stage, it includes 283K diverse and complex textual samples that feature explicit reasoning paths. For reinforcement learning with a verifiable reward, it contains an additional 31K complex textual examples and 21K multimodal questions, all of which are annotated with ground truths.
@@
As shown in Figure 2, the construction of GRAMMAR involves a multi-stage curation pipeline. We begin by amassing open-source reasoning datasets in Table 1, spanning various difficulty levels. This initial collection underwent rule-based filtering to ensure answer verifiability, excluding items like proof problems and those with difficult-to-verify ground truths. Subsequently, Qwen2.5-VL-7B-Instruct is employed for initial pruning of overly simple or complex questions. Then, Qwen2.5-VL-32B-Instruct is used to assess the remaining samples to classify them into ten difficulty levels. To maximize data diversity and minimize redundancy, we encoded questions using NV-Embedding-V2 (Lee et al., 2024), applied HDBSCAN (Campello et al., 2013) for clustering, assigned topics to clusters via Qwen2.5-7B-Instruct, and performed balanced sampling across both topics and difficulty strata.
@@ 4 Staged Reinforcement Optimization (SRO)
In Section 3, our data investigations and the curation of the GRAMMAR dataset highlight the necessity of high-quality, reasoning-focused data for developing advanced MLLM capabilities. In this section, we introduce Staged Reinforcement Optimization (SRO) to systematically cultivate robust reasoning and diverse competencies in MLLMs. Specifically, SRO contains two stages including multimodal RL and subsequently textual RL.
@@ 4.1 Stage 1: Multimodal RL
After the cold start training, the SRO framework commences with a dedicated Multimodal Reinforcement Learning (MRL) phase. This initial stage is pivotal for enabling the MLLM to ground textual concepts in visual information and execute cross-modal reasoning, primarily using the multimodal samples from our GRAMMAR dataset. We employ GRPO as the core RL algorithm for this phase. To ensure stable and effective learning, particularly when dealing with complex tasks and potentially sparse rewards common in multimodal settings, we propose a novel Prioritized Advantage Distillation (PAD) to improve gradient quality by addressing specific GRPO limitations.
@@ 4.1.1 Prioritized Advantage Distillation (PAD)
During multimodal RL training, we discover a significant challenge when applying GRPO in complex multimodal settings is “Gradient Stagnation”. This phenomenon refers to a reduction in learning efficacy due to a predominance of near-zero advantage estimates, which is particularly acute when dealing with sparse binary rewards. Essentially, if entire groups of generated responses yield uniform rewards (e.g., all correct or all incorrect), the resulting advantage signals become null, leading to zero policy gradients and thereby halting learning for those samples. This issue, also noted in concurrent works (Wang et al., 2025; Yu et al., 2025), can severely impede training progress. To specifically counteract gradient stagnation and enhance the efficiency of GRPO, we introduce Prioritized Advantage Distillation (PAD). PAD refines the training process by strategically focusing updates on the most informative samples within each batch, namely those exhibiting significant, non-zero advantage signals. The PAD mechanism, detailed below, operates on each batch after initial advantage estimation. It mainly contains the following three parts:
  • Per-Sequence Advantage Calculation: Compute the absolute advantage $|\hat{A}_{i}|$ for each sequence $i$ in the original batch $\mathcal{B}$, representing its learning signal magnitude.
  • Effective Sample Filtering: Form an “effective set” $\mathcal{E}$ by selecting sequences $i$ whose absolute advantage $|\hat{A}_{i}|$ falls within a specified informative range $T_{low},T_{high}]$. Critically, $T_{low}>0$ filters out stagnant (near-zero advantage) samples, ensuring that candidates for sub-sampling provide potentially useful learning signals.
@@
PAD thus directly counteracts gradient stagnation via a dual mechanism: first, by filtering out stagnant samples, and second, by prioritizing updates using informative, non-zero advantages from the remaining set. This selective optimization of the learning process ensures efficient computational resource allocation towards high-value samples. Consequently, PAD leads to enhanced training stability, improved learning efficiency, and more effective acquisition of complex reasoning skills.